OpenStudy (anonymous):

Given f(x) = sinx - cosx on [0, pi], on what interval(s) is f'(x) > 0?

OpenStudy (anonymous):

\[f'(x)>0 \]Means that \(f(x)\) is increasing.

OpenStudy (anonymous):

You could just take the derivative: \[ f'(x)=\cos(x)+\sin(x)>0 \]

OpenStudy (anonymous):



OpenStudy (anonymous):

So it looks like the interval is \([0,3\pi/4)\).

OpenStudy (anonymous):

so you just have to plug and chug in numbers? How do you do the problem with the math behind it?